f​ f(t) is the function that represents the temperature shown on the thermometer after t​ seconds, what is the closed interval for this​ application? Assume that the thermometer is at the starting temperature at time t=0.


enter your response here


​(Type your answer in interval​ notation.)



1
Expert's answer
2021-11-02T12:17:13-0400

It took 16 seconds for a mercury thermometer to rise from −20°C to 100°C when it was taken from a freezer and placed in boiling water.

If f(t) is the function that represents the temperature shown on the thermometer after t seconds, what is the closed interval for this application?

The closed intercal for this application is "[0, 16]."

Use the Mean Value Thorem

The function "f(t)" is continuous on "[0, 16]."

The function "f(t)" is differentiable on "(0, 16)."

Then there is a number "c" in "(0, 16)" such that


"f'(c)=\dfrac{f(16)-f(0)}{16-0}=\dfrac{100-(-20)}{160}=0.75 \degree C/s"